# Atomistic modeling of the channeling process with account for ionizing collisions: Preliminary results

### v1

@@ 3 Case studies: Preliminary results
$f_{\rm ch}(z)\approx{\rm erf}\left(\beta\sqrt{\varepsilon\over z}\right)$ (18)
where ${\rm erf}(.)$ stands for the error function.
The parameters $L_{\rm d}$ and $\beta$ can be defined by comparing the fitting formulae with the dependencies obtained in the simulations.
Figure 1: Channeling fractions $f_{{\rm ch},0}$ (dashed lines) and $f_{\rm ch}$ (solid lines) versus penetration distance $z$ for $\varepsilon=195,270$ and 855 MeV electrons in 200 $\mu$m thick diamond(111) crystal. Curves without symbols show the results obtained without accounting for the ionizing collisions; curves with symbols correspond to the dependencies calculated using the Molière potential to simulate the probability (7).
@@
The main conclusion, which one can draw comparing the dependencies presented in Fig. 1, is that for electrons the ionizing collisions can be disregarded in the quantitative description of the channeling motion. This can be understood recalling that negatively charged particles experience channeling motion in the vicinity of the atomic planes. Therefore, they experience rather close collisions (i.e. with comparatively small impact parameter) with the crystal atoms. At these impact parameters the contribution of the inelastic channels to the total cross section is much smaller than of the elastic collision, see, for example, [41].
Figure 2: Same as in Fig. 2 but for 270 and 500 MeV positrons in 400 $\mu$m thick diamond(110) crystal.
@@
The dependencies $f_{{\rm ch},0}$ (dashed lines) and $f_{\rm ch}$ (solid lines) obtained in the simulations of the 270 and 500 MeV positron channeling in a 400 $\mu$m thick single diamond(110) crystal performed without (no symbols) and with (open circles and squares) account for the ionizing collisions are shown in Fig. 2. In this case, at large penetration distances the values of both fractions calculated with account for the ionizing collisions are about 15 % (for $\varepsilon=270$ MeV) and 10 % (for 500 MeV) lower than those obtained ignoring the ionization. It is not surprising though that the impact of the ionizing collisions on the channeling efficiency of positrons is much stronger that for electrons. Positively charged particles effectively channel in the spatial regions between atomic planes, therefore, statistically they experience much more distant collisions with atoms than electrons. At large impact parameters the inelastic channels become dominant in the total cross section of the collision between an ultra-relativistic projectile and an atom. As a result, the increase in the scattering angle (and, consequently, in the transverse energy) occurs at higher rate leading to faster decrease in the channeling fractions with $z$.
4 Conclusions
We have reported the results of preliminary statistical analysis of the role which inelastic collisions with the crystal atoms play in the process of planar channeling of ultra-relativistic electrons and positrons of sub-GeV energies. The simulations have been performed by means of the MBN Explorer software package within the framework of relativistic classical molecular dynamics. The events of inelastic scattering of a projectile particle from crystal atoms have been incorporated into the classical framework according to their probabilities. On the basis of the simulations performed the dependence of the channeling fractions on the penetration distance has been obtained for several eelectron and positron beam energies.
More consistent analysis of the impact of the ionizing collisions on the dynamics of channeling as well as on the characteristics of the radiation emitted by channeling particles will be published elsehwere.

### v2

@@ III.1 Statistical analysis of the channeling process
$\widetilde{f_{\rm ch}}(z)\approx{\rm erf}\left(\beta\sqrt{\varepsilon\over z}\right)$ (20)
where ${\rm erf}(.)$ stands for the error function and $\beta$ is the fitting parameter. It is mentioned in the cited paper that if $\varepsilon$ is measured in MeV and $z$ in microns then as an initial guess for $\beta$ one can use the value of $(L_{\rm rad}U_{0})^{1/2}/106$ where $L_{\rm rad}$ is the radiation length (in cm) and $U_{0}$ is the depth of the interplanar potential (in eV) [19]. For sufficiently large distances ${\rm erf}\left(\beta\sqrt{\varepsilon/z}\right)\propto z^{-1/2}$.
To illustrate the applicability of (20) we present Fig. 5 (left graph), which compares dependencies $f_{\rm ch}(z)$ obtained from the simulations with the fitting function $\widetilde{f_{\rm ch}}(z)$.
@@
$\varepsilon=270$ MeV $\varepsilon=855$ MeV $\varepsilon=1.5$ GeV
${\cal A}$ $L_{\rm d}$ $L_{\rm p}$ $\beta$ ${\cal A}$ $L_{\rm d}$ $L_{\rm p}$ $\beta$ ${\cal A}$ $L_{\rm d}$ $L_{\rm p}$ $\beta$
C(110) 0.73 $4.4\pm 0.2$ $5.5\pm 0.3$ $0.104\pm 0.006$ 0.74 $10.5\pm 0.4$ $12.5\pm 0.5$ $0.092\pm 0.005$ 0.74 $16.3\pm 0.7$ $18.9\pm 0.9$ $0.087\pm 0.004$
0.74 $4.5\pm 0.1$ $5.7\pm 0.3$ 0.76 $12.0\pm 0.6$ $14.2\pm 0.7$ 0.76 $17.6\pm 0.7$ $20.8\pm 1.0$
C(111) 0.68 $7.0\pm 0.3$ $8.2\pm 0.4$ $0.125\pm 0.005$ 0.70 $17.0\pm 0.8$ $19.3\pm 1.0$ $0.111\pm 0.004$ 0.71 $27.4\pm 1.2$ $30.5\pm 1.5$ $0.110\pm 0.004$
0.69 $7.3\pm 0.2$ $8.6\pm 0.4$ 0.71 $18.6\pm 0.8$ $21.2\pm 1.1$ 0.72 $29.0\pm 1.2$ $32.7\pm 1.7$
Si(110) 0.64 $3.5\pm 0.2$ $5.0\pm 0.2$ $0.097\pm 0.005$ 0.66 $8.5\pm 0.4$ $11.4\pm 0.5$ $0.090\pm 0.004$ 0.74 $13.8\pm 0.6$ $17.9\pm 0.8$ $0.085\pm 0.003$
0.64 $3.7\pm 0.2$ $5.2\pm 0.2$ 0.67 $8.7\pm 0.4$ $11.8\pm 0.5$ 0.76 $14.1\pm 0.5$ $18.2\pm 0.7$
Si(111) 0.62 $5.1\pm 0.2$ $7.1\pm 0.3$ $0.111\pm 0.006$ 0.65 $13.2\pm 0.6$ $16.9\pm 0.8$ $0.102\pm 0.004$ 0.66 $20.3\pm 1.0$ $25.7\pm 1.2$ $0.096\pm 0.004$
0.62 $5.5\pm 0.3$ $7.4\pm 0.3$ 0.65 $13.6\pm 0.6$ $17.3\pm 0.8$ 0.67 $21.1\pm 1.2$ $26.5\pm 1.3$
Ge(110) 0.59 $5.7\pm 0.4$ $8.7\pm 0.3$ $0.057\pm 0.003$
0.59 $5.8\pm 0.2$ $8.8\pm 0.4$
Ge(111) 0.57 $7.9\pm 0.8$ $11.5\pm 0.5$ $0.061\pm 0.003$
0.58 $8.0\pm 0.8$ $11.6\pm 0.5$
Figure 4: Same as in Fig. 3 but for 530 MeV positrons.This paper presents a quantitative analysis of the impact of inelastic collisions of ultra-relativistic electrons and positrons, passing through oriented crystalline targets, on the channeling efficiency and on the intensity of the channeling radiation. The analysis is based on the numerical simulations of the channeling process performed using the MBN Explorer software package. The ionising collisions, being random, fast and local events, are incorporated into the classical molecular dynamics framework according to their probabilities. This methodology is outlined in the paper. The case studies presented refer to electrons with energy $\varepsilon$ ranging from 270 to 1500 MeV and positrons with $\varepsilon=530$ MeV incident on thick (up to 1 mm) single diamond, silicon and germanium crystals oriented along the (110) and (111) planar directions. In order to elucidate the role of the ionising collisions, the simulations were performed with and without account for the ionising collisions. The case studies presented demonstrate that both approaches yield highly similar results for the electrons. For the positrons, the ionising collisions reduce significantly the channeling efficiency. However, it has been observed that this effect does not result in a corresponding change in the radiation intensity. The dependencies for the (111) orientations are related to the ”wide” part of the channel.
Table 1 summarises the values of the aforementioned parameters namely the acceptance ${\cal A}$, the penetration distance $L_{\rm p}$, the asymptotic dechanneling length $L_{\rm d}$ and the fitting parameter $\beta$ calculated for $\varepsilon=270$, 855 and 1500 MeV electrons in different planar channels. For each channel shown, the results obtained with and without taking the ionising collisions into account are shown in the first and second rows, respectively.
Figure 5: Channeling fraction $f_{\rm ch}(z)$ for $1500$ MeV electrons (left) and 530 MeV positrons (right). The crystal orientation is indicated in the legends. The symbols show the results of the simulations with ionising collisions taken into account. The solid curves represent the fitting data calculated using Eq. (20) for electrons and the approximation $\widetilde{f_{\rm ch}}(z)\propto\exp(-z/L_{\rm r})$ for positrons. The values of the fitting parameters $\beta$ and $L_{\rm r}$ are given in Tables 1 and 2, respectively.
The results of the analysis of the channeling phenomenon for 530 MeV positrons are presented in Figs. 4 and 5 (right graph), and Table 2. In the figures, the curves shown for the (111) orientation of the crystals refer to the ”wide” part of the (111) channel, which exhibits much higher channeling efficiency than the ”narrow” part (see illustrative data in Appendix A.
As demonstrated in Fig. 4, the dependencies $f_{\rm ch,0}(z)$ and $f_{\rm ch}(z)$ exhibit two features that diverge distinctly from the electron channeling case.
@@
Firstly, it is evident that accounting for the ionising collisions results in a much steeper decrease of the fractions with the penetration distance. This is not surprising since positrons effectively channel in the spatial regions between atomic planes. Therefore, statistically, they experience much more distant collisions with atoms than channeling electrons. At large impact parameters, the inelastic channels dominate in the total cross section of the collision between an ultra-relativistic projectile and an atom. Consequently, the increase in the transverse momentum occurs at a higher rate, leading to a faster decrease of the fractions. It should be mentioned, nevertheless, that dechanneling due to the elastic collisions is not at all negligible. A comparison of the dependencies, calculated with and without ionisation losses, demonstrates that the elastic scattering provides from 30 % (in the case of diamond) up to 50 % (for Si(111) and Ge(110)) of the total decrease in the channeling fractions $f_{\rm ch,0}(z)$ and $f_{\rm ch}(z)$ at the crystal exit.
The second feature to be mentioned is that the replenishment of the channeling fraction due to the rechanneling phenomenon is much less efficient than in the case of electron channeling. Consequently, the dependencies $f_{\rm ch}(z)$ decrease much faster. In the range of penetration distances $z\leq 1000$ $\mu$m, considered in the simulations, the behaviour of these dependencies can be approximated as an exponentially decaying function $\propto\exp(-z/L_{\rm r})$, i.e. the same as for $f_{\rm ch,0}(z)$, Eq. (19), but with different fitting parameter $L_{\rm r}$. Figure 5right compares the dependencies $f_{\rm ch}(z)$ obtained from the simulations with the fitting function. The values of $L_{\rm r}$ are shown in Table 2 together with the acceptance ${\cal A}$ and the asymptotic dechanneling length $L_{\rm d}$.
Table 2: Acceptance ${\cal A}$, dechanneling length $L_{\rm d}$ and the parameter $L_{\rm r}$ (both in microns) calculated 530 MeV positrons with taking the ionising collisions into account. For the (111) orientation, the ${\cal A}$ and $L_{\rm d}$ values indicated in the upper row pertain to the ”wide” part of the channel, while the lower row corresponds to the ”narrow” part. The values of $L_{\rm r}$ refer to the ”wide” part only. The data refer to the positron beam of zero divergence.
C(110) C(111) Si(110) Si(111) Ge(110) Ge(111)
${\cal A}$ 0.94 0.78 0.93 0.75 0.92 0.74
0.17 0.18 0.17
$L_{\rm d}$ $305\pm 15$ $445\pm 15$ $640\pm 15$ $1060\pm 30$ $520\pm 10$ $900\pm 25$
$30\pm 3$ $44\pm 4$ $27\pm 2$
$L_{\rm r}$ $470\pm 10$ $620\pm 15$ $800\pm 20$ $1230\pm 25$ $615\pm 15$ $990\pm 20$
III.2 Spectral distribution of emitted radiation
In this Section, we present several case studies of the photon emission spectra by 855 and 1500 MeV electrons and 530 MeV positrons. A more extensive presentation and detailed analysis will be published elsewhere.
For each simulated trajectory, the spectral-angular distribution ${\rm d}^{3}E/{\rm d}(\hbar\omega)\,{\rm d}\Omega$ of the radiation of energy $\hbar\omega$ emitted within the solid angle ${\rm d}\Omega\approx\theta{\rm d}\theta{\rm d}\phi$ (where $\theta$ and $\phi$ are the polar angles of the emission direction) is calculated numerically following the algorithm outlined in Ref. [15]. The spectral distribution of energy radiated within the cone $\theta_{0}\ll 1$ along the incident beam and averaged over all trajectories is calculated using the following formula:
${{\rm d}E(\theta\leq\theta_{0})\over{\rm d}(\hbar\omega)}={1\over N_{0}}\sum_{n=1}^{N_{0}}\int\limits_{0}^{2\pi}{\rm d}\phi\int\limits_{0}^{\theta_{0}}\theta{\rm d}\theta\,{{\rm d}^{3}E_{n}\over{\rm d}(\hbar\omega)\,{\rm d}\Omega}.$ (21)
The sum is carried out over all simulated trajectories of the total number $N_{0}$; as a result it takes into account the radiation emitted from the channeling segments of the trajectories as well as from those corresponding to the non-channeling regime.
Figure 6: Spectral distribution of the radiation emitted by 855 MeV (upper row) and 1500 MeV (lower row) electrons channeled in (110) and (111) channels in the diamond, silicon and germanium crystals, as indicated. The solid curves show the results obtained taking into account the ionising collisions (labelled ’yes’ in the common legends placed in the upper right graph). The dashed curves represent the dependencies calculated without ionising collisions (labelled ’no’ in the legend). The spectra shown correspond to the crystal thicknesses $L=200$ and 400 $\mu$m and the emission cones $\theta_{0}=3$ and 1 mrad for the 855 and 1500 MeV projectiles, respectively. The dashed-dotted lines show the spectrum of the incoherent bremsstrahlung emitted in the amorphous target of the same thickness.
The spectral distributions of radiation emitted by electrons are presented in Fig. 6. The upper row corresponds to 855 MeV electrons propagating through crystalline targets with a thickness of 200 $\mu$m, while the distributions for 1500 MeV electrons in 400 $\mu$m thick crystals are shown in the graphs in the lower row. For the lower energy, the spectra were computed for the cone with $\theta_{0}=3$ mrad, which exceeds the natural emission cone $1/\gamma$ by a factor of approximately five. For the higher energy, the cone considered 3 mrad is approximately $3/\gamma$. The common legend placed in the top-right graph indicates the crystals’ orientation as well as marks the distributions calculated with account for ionising collisions (’yes’) and without the ionising collisions (’no’). For the sake of comparison, the dashed-dotted line in each graph shows the spectrum of bremsstrahlung emitted in the amorphous medium of the same thickness. These spectra were calculated within the Bethe–Heitler approximation [42, 43].
For both energies of the electron beam the distributions are calculated in the photon energy ranges where the spectrum is dominated by the channeling radiation [44]. For the purpose of this paper it is important to note that, similar to the channeling fractions, the account for the ionising collision does not affect the spectral distributions. The most significant discrepancy (of approximately 10 per cent in the maximum of the spectra) between the ’yes’ and ’no’ curves is observed for 855 MeV electrons in the diamond targets.
Figure 7: Same as in Fig. 6 but for 530 MeV positrons. In all cases the crystal thicknesses is $L=1000$ $\mu$m and the emission cone $\theta_{0}=1$ mrad. The dashed-dotted lines show the spectrum of the incoherent bremsstrahlung calculated in the Bethe-Heitler approximation.
The results for the spectral distributions of radiation by 530 MeV positrons in $L=1000$ $\mu$m-thick diamond, silicon and germanium oriented crystals are presented in Fig. 7. In all cases the emission cone is $\theta_{0}=1\,\mbox{mrad}\approx 3/\gamma$.
Considering a significant decline in channeling efficiency due to ionising collisions, see Fig. 4, one could expect to have a similar impact on the spectral distribution of the channeling radiation. It is seen that when ionising collisions are taken into account (the solid curves marked ’yes’ in the legend), the peak value of the channeling radiation decreases by approximately 20-25 per cent for diamond and germanium, and remains virtually unchanged for silicon crystal. A possible explanation for this feature is based on the assumption that, although ionising collisions lead to a reduction in the average length of the channeling segments they also lead to a gradual increase in the amplitude of the channeling oscillations for the channeling particle. The former phenomenon tends to reduce the intensity of the channeling radiation, while the latter has the opposite effect. A more detailed quantitative analysis of this assumption is currently on the way.
IV Conclusions
A quantitative analysis of the impact of inelastic collisions on the channeling efficiency and the radiation emission by 270-1500 MeV electrons and 530 MeV positrons incident on diamond, silicon and germanium single crystals oriented along the (110) and (111) planar directions has been reported. The crystals thickness along the incident beam was 200-400 $\mu$m for the electrons and 1000 $\mu$m for the positrons.
The simulations were performed by means of the MBN Explorer software package within the framework of classical relativistic MD. The incorporation of inelastic scattering events of projectile particles from crystal atoms into the classical framework was conducted according to their probabilities.
In order to elucidate the role of the ionising collisions, calculations were performed of the channeling fractions versus the penetration distance and of the spectral distribution of the emitted radiation, with and without account for these collisions.
The case studies presented for the electron beams have demonstrated that both approaches lead to very close results. The largest impact of ionising collisions on the values of the channeling fractions and of the channeling radiation intensity, seen for the diamond targets, is on the level of ten per cent. This discrepancy is significantly reduced in the heavier crystals. A comparatively weak role for ionising collisions can be explained by recalling that electrons channel in the vicinity of atomic planes. In this regime, the change in the transverse energy of the particle is mainly due to elastic scattering from the static atomic potential rather than inelastic scattering from atomic electrons.
The situation is less straightforward for positrons. They channel into the spatial regions between the atomic levels and therefore experience more distant collisions in which the inelastic channels dominate in the total cross section of a scattering from an atom. The simulations of the channeling motion have supported this conclusion leading to a much faster decrease in the channeling fractions with increasing penetration distance. The dechanneling rate due to inelastic collisions increases, ranging from 50 per cent (for Si(111) and Ge(110)) up to 70 per cent (for diamond) of the total dechanneling rate. However, the increase in the dechanneling rate does not lead to a comparable change in the radiation emission spectra. This phenomenon may be explained by the assumption that ionising collisions reduce the average length of the channeling segments, but increase the amplitude of the channeling oscillations. The former reduces the intensity of the channeling radiation, while the latter increases it. A more detailed quantitative analysis of the particle dynamics and a more comprehensive presentation of the data on the radiation emission spectra will be presented elsewhere.
@@ Acknowledgements
Appendix A The (111) interplanar potentials in diamond, silicon and germanium crystals
The results presented in the main text have been obtained using the all-atom relativistic MD approach implemented in the MBN Explorer software package [14, 15, 17, 15]. This approach goes beyond the continuous potential model [1]. Therefore, the numerical data presented below in this section are for illustrative purposes only.
Figure 8: Continuous interplanar potential for a positron in the diamond (C), silicon and germanium crystals with the (111) orientation. The vertical dashed lines mark the positions of the crystallographic (111) planes and illustrate the wide and narrow channels. The solid and dashed curves represent the potentials calculated using the Molière (M) [30] and Fernandes Pacios (F-P) [31] parametrisations of the atomic potentials, respectively.
In crystals of the diamond group (diamond itself, silicon, germanium) the (111) crystallographic planes are arranged with two different alternating spacings. Expressed in terms of the lattice constant a, the wide interplanar distance is $d_{\rm W}=0.75a/\sqrt{3}$ and the narrow spacing is three times less, $d_{\rm N}=0.25a/\sqrt{3}$. The channeling motion of a positively charged projectile particle occurs in between two neighbouring planes. Consequently, in the crystal with the (111) orientation there are wide and narrow positron channels.
Figure 8 illustrates the continuous interplanar potentials in the wide and narrow (111) positron channels in the diamond, silicon and germanium crystals. The vertical dashed lines mark the positions of the crystallographic (111) planes. In each case, the interplanar potential is calculated by summing the continuous potentials U of individual planes. In turn, the potentials U are calculated by averaging out the potentials of individual atoms, considering their displacement from the nodal positions due to the thermal vibrations. The dependencies presented refer to a temperature $T=300$ C. The solid and dashed curves in the figure represent the potentials calculated using the Molière (M) [30] and Fernandes Pacios (F-P) [31] parametrisations of the atomic potentials, respectively. More technical details on the calculation of the continuous planar potentials can be found in Ref. [19].
It is clear that, in all cases, the potential well in the wide channel is much deeper. This, and the big difference in the widths of the channels, indicates that the positron channeling efficiency of the wide channels is much higher than that of the narrow channels. This is also shown by the numerical simulations, which show that the dechanneling length, $L_{\rm d}$ in narrow channels is one order of magnitude less than in wide channels (see Table 2).